# CUB: Benchmarking Context Utilisation Techniques for Language Models

### v1

@@ 5 Features Impacting Context Utilisation
To deepen our understanding of the results on CUB, we complement the benchmark with an analysis of features likely to impact context utilisation. Our goal is to better understand why certain CMTs and LMs work well or not. We study features on a model and input level, described below.
@@ 5.1 Model Features
By virtue of the large LM coverage in CUB, we are able to measure multiple salient model features. We analyse model size, whether the model is instruction-tuned and strength of model memory. To control for external confounders related to model family and implementation, we only measure correlations with model size and instruction-tuning across Qwen models. Strength of model memory is measured as the softmaxed logits for the top token predicted by the LM when only the query is provided (without context).
@@ 5.2 Input Features
We measure multiple input characteristics found to impact context utilisation for humans and/or LMs. By considering context length and Flesch reading ease score, we aim to measure whether the context is difficult to understand (Gao et al., 2024; Vladika and Matthes, 2023). Using distractor rate, we aim to measure whether the context contains distracting information (Shaier et al., 2024). With query-context overlap we also aim to measure query-context similarity (Wan et al., 2024). Lastly, we check the answer position (Liu et al., 2024) and if the evaluated LMs find the context relevant. More details on the detection of the features can be found in Appendix G.
@@ 5.3 Metric for Feature Impact
By virtue of the unified setup of CUB, we can study correlation coefficients to investigate the impact of different input and model features with a low risk of confounders. We use Spearman’s $\rho$ to measure the impact of features on context utilisation, proxied by BCU.
@@ 6 Main Results on CUB
Figure 3: $\mathrm{BCU}$ scores for the evaluated context utilisation manipulation methods applied to the evaluated models and datasets. ‘Total’ denotes the averaged performance across all context types. A high $\mathrm{BCU}$ score is desirable regardless of context type.
The CUB results can be found in Figures 3 and 4, $\mathrm{CCU}$ scores and more detailed results can be found in Appendix A. We structure the results analysis around a set of main findings.
@@ 6.1 Overall Trends
We first note that the $\mathrm{BCU}$ and $\mathrm{CCU}$ scores in Figures 3 and 5, respectively, support the same trends and focus the analysis on the $\mathrm{BCU}$ results.
Context utilisation improves with model size.
From Figure 3, we note how larger Regular LMs generally outperform smaller LMs when all context types are taken into consideration for NQ and DRUID. On NQ, the best performing model is Qwen 32B, and on DRUID the best performing model is Command A. Notably, applying a CMT to a small LM can lead to context utilisation on par with that of a regular larger LM, such as Fine-tuning Qwen 7B compared to Regular Qwen 32B on NQ. Meanwhile, on CounterFact, we observe how Regular model performance across all contexts generally decreases when model size is increased. This is counter-intuitive and we attribute the phenomena to the artificial nature of the dataset, which likely confuses the larger LMs. In addition, we know the NQ and DRUID datasets to be more difficult, demanding greater model capacity. This shows how it is insufficient to evaluate context utilisation only on simple datasets like CounterFact.
Most CMTs show an inflated performance on conflicting CounterFact contexts.
All LMs that do not already have a perfect BCU score on the conflicting CounterFact contexts improve to a perfect score of 1.0 under Prompting, PH3 +context, and Fine-tuning. However, similar improvements cannot be observed for the same CMTs on NQ or DRUID. These results show how CMTs proven to work well in simpler settings are not guaranteed to work equally well in more complex settings, proving the necessity of holistic tests. A deeper analysis of the inflated CMT performance on CounterFact is provided in Appendix A.
6.2 CMT Comparsion
@@
There is a conflict between optimising for utilisation of relevant contexts and robustness to irrelevant contexts.
As each CMT exhibits trade-offs across context types or only marginal differences from Regular, the overall CMT $\Delta$ values (Total) converge to near zero across NQ and DRUID. Consequently, we find no CMT that is superior. For instance, PH3 +context shows consistent improvements over Regular in conflicting contexts, but underperforms when applied to irrelevant contexts. Conversely, ACD, which handles irrelevant context effectively, performs worse in the conflicting context setting. Unsurprisingly, these findings highlight that the effectiveness of each CMT is closely tied to the alignment between the objective of the CMT and the type of context being provided. RAG practitioners knowing beforehand that their retrieval system is e.g. prone to return irrelevant information, may prioritise robustness over strong context utilisation and can select e.g. ACD as the CMT most suitable to their needs.
Prompting-based CMTs, such as Prompting and Multi-agent, show relatively stable performance across context types, without substantial drops in $\Delta$. Compared to other CMTs, they offer this robustness with lower optimisation and implementation costs. Multi-agent shows clear gains in irrelevant contexts but limited efficacy in gold and conflicting settings. This suggests that LMs are capable of identifying irrelevant contexts, but remain limited in effectively utilising relevant ones.
In realistic RAG scenarios, it will not be known beforehand what types of context will be provided to the LM. Therefore, it is important that CMTs work optimally across all context types. Our work shows that while we have CMTs that work well for relevant or irrelevant contexts alone, there currently are no CMTs that handle both relevant and irrelevant contexts well.
@@ 6.3 Impact of Model and Input Features
See Tables 6 and 7 for Spearman’s $\rho$ between BCU and the features described in Section 5. Results are averaged across models.
@@
Corroborating our findings in Section 6.1, we observe a positive correlation with model size ($\rho\approx 0.3$) on DRUID gold contexts. Multi-agent also works significantly better with bigger LMs on DRUID gold contexts ($\rho=0.42$). In addition, we observe a positive correlation with model size on NQ gold contexts ($\rho\in[0.20,0.37]$). For CounterFact, we observe how model size does not correlate with performance.
@@
Conversely, instruction-tuning is clearly detrimental for conflicting CounterFact contexts ($\rho\leq-0.36$), potentially because the LMs have been more tuned to be critical of unreliable information, as opposed to following a pure causal language modelling objective.
@@
We measure low correlation values (below 0.3) across all settings for answer position in the context and Flesch reading ease score, and have thus omitted them in Table 7. Previous work has already found the Flesch reading ease score to show low correlations with LM context utilisation; our work further supports this finding (Hagström et al., 2024). Liu et al. (2024) found the answer position impactful for the utilisation of long contexts. CUB does not contain equally long contexts, which potentially explains why we do not see the same impact of answer position.
@@ 7 Conclusion
We introduce CUB, a benchmark that evaluates CMTs across diverse context types, datasets, and models. Under CUB, we evaluate a representative set of CMTs, covering varying context utilisation objectives and techniques. Results on CUB reveal a trade-off across most CMTs between robustness to irrelevant context and faithful utilisation of relevant context. Our analysis of features impacting context utilisation highlights the strong influence of model features, while input features have limited impact when analysed in separation . Overall, our findings highlight the need for holistic testing, as tests on synthesised datasets may show inflated performance, and the need for CMTs that can adapt to varied context conditions. Taken together, our work paves the way for the development of more effective RAG systems.
@@
CUB only incorporates contexts with lengths of up to that of a paragraph. It would also be relevant to evaluate CMTs in long-context settings. The long-context setting was not included in CUB, and left for future work, as it is fundamentally different from the normal context setting studied in CUB, posing new challenges for context utilisation and its evaluation, associated with a different set of CMTs (Shaham et al., 2023; Zhang et al., 2024a; Min et al., 2023; Zhang et al., 2024b).
While the dataset selection for CUB was performed to cover a wide span of task difficulty and RAG scenarios, the insights provided by CUB are limited to those derived from the underlying datasets. Moreover, all datasets are in English, leaving open the question of whether the findings generalise across languages (Chirkova et al., 2024).
Lastly, CUB does not explicitly consider datasets involving temporal dynamics, while it would be interesting to study. Time-sensitive information may lead to naturally occurring conflicts in context, adding nuance to the analysis of context utilisation (Loureiro et al., 2022; Xiong et al., 2024).
@@ Acknowledgments
[Uncaptioned image][Uncaptioned $\end{array}$ This research was co-funded by the European Union (ERC, ExplainYourself, 101077481), by the Pioneer Centre for AI, DNRF grant number P1, as well as by The Villum Synergy Programme. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them.
@@ A.1 CUB results
The exact CUB results can be found in Tables 3 and 4. $\mathrm{CCU}$ scores can be found in Figure 5. For the $\mathrm{CCU}$ scores, we note that they generally follow the same trends as the $\mathrm{BCU}$ scores in Figure 3; some CMTs perform better on gold, conflicting or irrelevant contexts, while none are superior when all context types are taken into consideration. The only disparate trend at odds with the $\mathrm{BCU}$ scores is that Fine-tuning Qwen models that have been instruction-tuned stand out by performing extra poorly with respect to $\mathrm{CCU}$ score. We hypothesise that this is a consequence of an increase in $P_{M}(t_{C}|Q)$ (i.e. prediction probability without context) from the fine-tuning, yielding less room for improvement in prediction confidence when context is introduced.
@@ A.2 Analysis of inflated CMT performance on CounterFact
The inflated performance on CounterFact, observed in Figures 3 and 4, can potentially be explained by a suboptimal default prompt for CounterFact. Following previous work, the default prompt only contained the example to be completed, without any additional instructions or few-shot examples. For NQ and DRUID, the default prompt contained task instructions and few-shot examples. Furthermore, we observe how Prompting performs best on CounterFact on average, with a near perfect performance, indicating that a better default prompt may have neutralised any additional improvements from other CMTs. This raises the question of whether certain CMTs only address low context utilisation when caused by poor prompting, finding no leverage if the prompt already is adequate.
@@ A.4 Performance of Relevance Judgement
For the Multi-agent technique, we investigate whether instruction-tuned LMs are capable of identifying irrelevant context when explicitly prompted to do so. According to Table 5, the Multi-agent approach demonstrates strong performance in detecting irrelevant contexts and in recognising gold contexts as relevant. Although it does not reliably maintain a closed-book response when directly generating responses (i.e. Regular), it can accurately detect irrelevance when equipped with an explicit relevance assessment setup.
@@ A.5 Features Impacting Context Utilisation
See Table 6 for the correlation values between model features and context utilisation. See Table 7 for the correlation values between input features and context utilisation.
Dataset Context CMT Corr.
Model size
DRUID Gold Multi-agent 0.42
DRUID Gold ACD 0.41
NQ Gold PH3 +memory 0.37
DRUID Gold Regular 0.36
DRUID Gold Prompting 0.36
NQ Conflicting PH3 +memory 0.33
NQ Gold Regular 0.20
NQ Irrelevant Regular 0.14
NQ Conflicting Regular 0.09
CounterFact Gold Regular 0.04
CounterFact Irrelevant Regular 0.02
CounterFact Conflicting Regular -0.01
DRUID Conflicting Regular -0.08
DRUID Irrelevant Regular -0.20
DRUID Irrelevant PH3 +memory -0.33
CounterFact Conflicting Fine-tuning -0.33
DRUID Irrelevant COIECD -0.44
Instruct tuned
DRUID Conflicting PH3 +memory 0.77
DRUID Irrelevant PH3 +context 0.65
DRUID Conflicting ACD 0.54
DRUID Conflicting Prompting 0.46
DRUID Conflicting Regular 0.40
DRUID Conflicting COIECD 0.34
DRUID Irrelevant Regular 0.29
NQ Gold Regular 0.13
CounterFact Irrelevant Regular 0.12
NQ Irrelevant Regular 0.06
NQ Conflicting Regular 0.05
CounterFact Gold Regular 0.01
DRUID Gold Regular -0.19
CounterFact Conflicting Regular -0.36
DRUID Gold ACD -0.38
CounterFact Conflicting PH3 +context -0.43
DRUID Gold PH3 +memory -0.72
Strength of memory
DRUID Conflicting PH3 +memory 0.54
NQ Irrelevant Fine-tuning 0.47
NQ Irrelevant ACD 0.39
CounterFact Irrelevant Fine-tuning 0.39
NQ Irrelevant Prompting 0.39
NQ Irrelevant COIECD 0.38
DRUID Conflicting ACD 0.37
NQ Irrelevant Regular 0.37
CounterFact Irrelevant Regular 0.35
DRUID Conflicting Prompting 0.34
CounterFact Irrelevant ACD 0.32
CounterFact Irrelevant PH3 +memory 0.31
CounterFact Irrelevant COIECD 0.30
DRUID Conflicting Regular 0.26
NQ Gold Regular 0.18
DRUID Irrelevant Regular 0.15
NQ Conflicting Regular 0.09
CounterFact Gold Regular 0.04
DRUID Gold Regular 0.02
CounterFact Conflicting ACD -0.31
CounterFact Conflicting COIECD -0.42
DRUID Gold PH3 +memory -0.43
CounterFact Conflicting Regular -0.44
Table 6: Spearman’s $\rho$ between BCU and different model aspects. Correlation values for Regular or with an absolute value above 0.3 are shown. Correlation values with an absolute value below 0.3 are marked in gray. Significant correlation values (p-value < 0.05) are marked in bold.
@@
Table 7: Spearman’s $\rho$ between BCU and different input aspects. Correlation vallues for Regular or with an absolute value above 0.3 are shown. Correlation values with an absolute value below 0.3 are marked in gray. Significant correlation values (p-value < 0.05) are marked in bold.

### v2

@@ 5 Features Impacting Context Utilisation
To deepen our understanding of the results on CUB, we complement the benchmark with a sophisticated analysis of features likely to impact context utilisation. Our goal is to uncover the fundamental factors that determine why certain CMTs and LMs achieve superior performance. We systematically investigate features at both model and input levels, providing unprecedented insights into the mechanisms underlying context utilisation.
@@ 5.1 Model Features
By virtue of the unprecedented LM coverage in CUB, we can systematically measure multiple salient model features. We analyse model size, whether the model is instruction-tuned, and strength of model memory. To control for external confounders related to model family and implementation differences, we carefully measure correlations with model size and instruction-tuning exclusively across Qwen models. Strength of model memory is quantified as the softmaxed logits for the top token predicted by the LM when only the query is provided (without context).
Figure 2: $\mathrm{BCU}$ scores for the evaluated context utilisation manipulation methods applied to the evaluated models and datasets. ‘Total’ denotes the averaged performance across all context types.
@@ 5.2 Input Features
We measure multiple input characteristics found to impact context utilisation for humans and/or LMs (see Appendix˜H). By considering context length and Flesch reading ease score, we aim to measure whether the context is difficult to understand (Gao et al., 2024; Vladika and Matthes, 2023). Using distractor rate, we aim to measure whether the context contains distracting information (Shaier et al., 2024). With query-context overlap we also aim to measure query-context similarity (Wan et al., 2024). Lastly, we check the answer position (Liu et al., 2024) and if the evaluated LMs find the context relevant.
@@ 5.3 Metric for Feature Impact
By virtue of the unified and controlled setup of CUB, we can systematically study correlation coefficients to investigate the impact of different input and model features with minimal risk of confounding factors. We employ Spearman’s $\rho$ to measure the impact of features on context utilisation, proxied by BCU, providing robust statistical insights into the factors that drive CMT effectiveness.
@@ 6 Main Results on CUB
The CUB results can be found in Figures˜2 and 3; $\mathrm{CCU}$ scores together with additional results and analyses can be found in Appendix˜A. We structure our analysis around a set of key findings about CMT effectiveness.
@@ 6.1 Overall Trends
We first note that the $\mathrm{BCU}$ and $\mathrm{CCU}$ scores in Figures˜2 and 4, respectively, support the same trends and focus the analysis on the $\mathrm{BCU}$ results.
Context utilisation scales with model size – but with surprising exceptions.
From Figure˜2, we observe that larger Regular LMs generally achieve superior performance across all context types for NQ and DRUID, following expected scaling laws. On NQ, the best performing model is Qwen 32B, while on DRUID, Command A leads performance. Remarkably, our results reveal that applying the right CMT to a smaller LM can achieve context utilisation performance on par with much larger regular LMs – for instance, Fine-tuning applied to Qwen 7B matches the performance of Regular Qwen 32B on NQ. However, CounterFact reveals a striking counter-intuitive pattern: Regular model performance across all contexts actually decreases as model size increases. We attribute this unexpected phenomenon to the artificial nature of the dataset, which appears to confuse larger LMs. This finding, combined with the fact that NQ and DRUID datasets demand greater model capacity, underscores a critical insight: evaluating context utilisation solely on simple datasets like CounterFact provides a fundamentally misleading picture of CMT effectiveness.
Most CMTs exhibit inflated performance on conflicting CounterFact contexts.
We observe a striking phenomenon: all LMs that do not already achieve perfect BCU scores on conflicting CounterFact contexts leap to a perfect score of 1.0 under Prompting, PH3 +context, and Fine-tuning. However, these same CMTs fail to deliver comparable improvements on the more realistic NQ or DRUID datasets. These results expose a critical flaw in current evaluation practices – CMTs that appear highly effective in simplified settings may fail to generalise to real-world complexity. This finding demonstrates the essential need for holistic evaluation frameworks like CUB.
6.2 CMT Comparison
@@
We uncover a fundamental trade-off between sensitivity to relevant contexts versus robustness to irrelevant contexts.
Our analysis reveals a striking pattern: each CMT exhibits inherent trade-offs across context types, with overall $\Delta$ values (Total) converging to near zero across NQ and DRUID. This convergence exposes that no single CMT emerges as universally superior. For instance, PH3 +context demonstrates consistent improvements over Regular in conflicting contexts, but significantly underperforms when applied to irrelevant contexts. Conversely, ACD, while effectively handling irrelevant contexts, shows degraded performance in conflicting context scenarios.
In realistic RAG scenarios, the type of context provided to the LM cannot be predicted in advance. Therefore, the ideal CMT must perform optimally across all context types. Our comprehensive evaluation reveals that while we have identified CMTs that excel with either relevant or irrelevant contexts in isolation, no existing CMT can robustly handle both relevant and irrelevant contexts simultaneously. This represents a fundamental gap in current CMT capabilities that needs to be addressed.
Meanwhile, prompting-based CMTs, such as Prompting and Multi-agent, demonstrate remarkably stable performance across context types, avoiding the dramatic performance drops observed with other approaches. Compared to more complex CMTs, they offer this robustness with lower optimisation and implementation costs, making them attractive for practical deployment. Multi-agent shows particularly promising results with clear gains in irrelevant contexts, though its efficacy remains limited in gold and conflicting settings. This pattern suggests that while LMs demonstrate competence in identifying irrelevant contexts, they face fundamental limitations in effectively utilising relevant information.
@@ 6.3 Impact of Model and Input Features
Dataset Context CMT Corr.
Model size
DRUID Gold Multi-agent 0.42
DRUID Gold ACD 0.41
NQ Gold PH3 +memory 0.37
DRUID Gold Regular 0.36
DRUID Gold Prompting 0.36
NQ Conflicting PH3 +memory 0.33
NQ Gold Regular 0.20
NQ Irrelevant Regular 0.14
NQ Conflicting Regular 0.09
CounterFact Gold Regular 0.04
CounterFact Irrelevant Regular 0.02
CounterFact Conflicting Regular -0.01
DRUID Conflicting Regular -0.08
DRUID Irrelevant Regular -0.20
DRUID Irrelevant PH3 +memory -0.33
CounterFact Conflicting Fine-tuning -0.33
DRUID Irrelevant COIECD -0.44
Instruct tuned
DRUID Conflicting PH3 +memory 0.77
DRUID Irrelevant PH3 +context 0.65
DRUID Conflicting ACD 0.54
DRUID Conflicting Prompting 0.46
DRUID Conflicting Regular 0.40
DRUID Conflicting COIECD 0.34
DRUID Irrelevant Regular 0.29
NQ Gold Regular 0.13
CounterFact Irrelevant Regular 0.12
NQ Irrelevant Regular 0.06
NQ Conflicting Regular 0.05
CounterFact Gold Regular 0.01
DRUID Gold Regular -0.19
CounterFact Conflicting Regular -0.36
DRUID Gold ACD -0.38
CounterFact Conflicting PH3 +context -0.43
DRUID Gold PH3 +memory -0.72
Strength of memory
DRUID Conflicting PH3 +memory 0.54
NQ Irrelevant Fine-tuning 0.47
NQ Irrelevant ACD 0.39
CounterFact Irrelevant Fine-tuning 0.39
NQ Irrelevant Prompting 0.39
NQ Irrelevant COIECD 0.38
DRUID Conflicting ACD 0.37
NQ Irrelevant Regular 0.37
CounterFact Irrelevant Regular 0.35
DRUID Conflicting Prompting 0.34
CounterFact Irrelevant ACD 0.32
CounterFact Irrelevant PH3 +memory 0.31
CounterFact Irrelevant COIECD 0.30
DRUID Conflicting Regular 0.26
NQ Gold Regular 0.18
DRUID Irrelevant Regular 0.15
NQ Conflicting Regular 0.09
CounterFact Gold Regular 0.04
DRUID Gold Regular 0.02
CounterFact Conflicting ACD -0.31
CounterFact Conflicting COIECD -0.42
DRUID Gold PH3 +memory -0.43
CounterFact Conflicting Regular -0.44
Table 3: Spearman’s $\rho$ between BCU and model features. Correlation values for Regular or with an absolute value above 0.3 are shown. Correlation values with an absolute value below 0.3 are marked in gray. Significant correlation values (p-value < 0.05) are marked in bold. Results are measured across models.
See Tables˜3 and 7 for Spearman’s $\rho$ between BCU and the features described in Section˜5. Our findings are as follows.
@@
Corroborating our findings in Section˜6.1, we observe a positive correlation with model size ($\rho\approx 0.3$) on DRUID gold contexts. Multi-agent also works significantly better with bigger LMs on DRUID gold contexts ($\rho=0.42$). In addition, we observe a positive correlation with model size on NQ gold contexts ($\rho\in[0.20,0.37]$).
@@
Conversely, instruction-tuning is clearly detrimental for conflicting CounterFact contexts ($\rho\leq-0.36$), potentially because this makes the LMs more critical of unreliable information, as opposed to performing pure causal language modelling.
@@
We measure low correlation values (below 0.3) across all settings for answer position in the context and Flesch reading ease score, and have thus omitted them in Table˜7. Previous work has already found the Flesch reading ease score to show low correlations with LM context utilisation; our work further supports this finding (Hagström et al., 2024). Liu et al. (2024) found the answer position impactful for the utilisation of long contexts. CUB does not contain equally long contexts, which potentially explains why we do not see the same impact of answer position.
@@ 7 Conclusion
We present CUB, the first comprehensive benchmark for systematically evaluating CMTs across diverse contexts, datasets, and model architectures. Using CUB, we conduct the most extensive comparison to date of seven representative CMTs, revealing a core trade-off between robustness to irrelevant context and effective use of relevant information. Our analysis shows that model characteristics play a dominant role in context utilisation, while input features alone are weak predictors. Crucially, we uncover major flaws in current evaluation practices – many CMTs perform well on synthetic datasets but falter in realistic settings. These findings underscore the need for holistic evaluation frameworks and set a new standard for developing context manipulation techniques in real-world RAG environments.
@@
While context can be characterised by various dimensions, including its length or the number of documents provided, our current study specifically focuses on context type. We view the investigation of longer or multi-document contexts as important fields for further research, yet orthogonal to our current focus, as explained below.
Long-context settings (e.g., over 4k tokens) have recently garnered significant attention (Liu et al., 2024), and evaluating CMTs in these scenarios is an important direction for future research. However, this paper focuses on a different set of challenges – those associated with standard-length contexts – making long-context evaluation out of scope. Long-context tasks involve distinct issues in context utilisation and require different evaluation strategies and CMTs (Shaham et al., 2023; Zhang et al., 2024a; Min et al., 2023; Zhang et al., 2024b). Our goal with CUB was to introduce a foundational benchmark that enables rigorous comparisons between CMTs across diverse types of contexts and domains. Starting with standard-length contexts was a natural choice, as they remain both prevalent and crucial. Even within this scope, CUB uncovers valuable insights into CMT behavior. Future work can extend on the foundation provided by CUB to include long-context benchmarks.
The multi-document setting – where multiple contexts are provided to the model – is another important area for future research. However, CUB focuses on single-document settings to allow controlled and meaningful evaluation of diverse CMTs across different context types. Evaluating CMTs in a multi-document setup introduces significant complexity, as it becomes challenging to disentangle the influence of individual documents from their combined effects. Furthermore, many of the recent and widely used CMTs included in CUB were designed specifically for single-document scenarios and are not yet compatible with multi-document inputs. As such, CUB provides a necessary and robust foundation, which future work can build upon to address the more intricate challenges of multi-document settings.
CUB is built on a carefully curated selection of nine models, chosen to balance recency, diversity in model size, instruction tuning, and relevance to RAG-specific applications. While a limitation of CUB is its coverage of only a subset of available models, the thoughtful selection ensures that the insights it provides are broadly informative and likely to generalise. Additionally, CUB is designed to be extensible, allowing future inclusion of a wider range of models.
While the CMTs included in CUB represent the current state of the art and cover all major categories, the rapidly evolving landscape means that many promising methods remain to be benchmarked. However, by open-sourcing CUB, we provide the research community with a robust foundation for systematic evaluation of both existing and future CMTs.
The datasets used in CUB were carefully selected to span a range of task difficulties and reflect common RAG scenarios. However, the insights derived from CUB are inherently constrained by the characteristics of these underlying datasets. Additionally, since all datasets are in English, it remains an open question whether the findings generalise to other languages (Chirkova et al., 2024). Future work could extend CUB to cover more datasets as well as non-English datasets.
Finally, CUB does not explicitly address datasets involving temporal dynamics, which represent an interesting avenue for future research. Time-sensitive information can introduce natural conflicts within the context, offering a richer setting to analyse context utilisation (Loureiro et al., 2022; Xiong et al., 2024). Additionally, while the context types included in CUB were carefully selected based on prior research, it is possible that other context types have been overlooked.
@@ Acknowledgments
[…]
@@ A.1 CUB results
The exact BCU scores on CUB can be found in Table˜4. We also report the accuracy scores on CUB in Table˜5. For CounterFact and DRUID, accuracy is measured based on whether the first generated token is the same as the first gold token. For NQ, for which the correct answer may be different permutations of the same set of tokens, we measure accuracy based on whether the first output token (e.g. “July”) matches any of the tokens in the answer (e.g. “15 July”).
$\mathrm{CCU}$ scores can be found in Figure˜4. For the $\mathrm{CCU}$ scores, we note that they generally follow the same trends as the $\mathrm{BCU}$ scores in Figure˜2; some CMTs perform better on gold, conflicting or irrelevant contexts, while none are superior when all context types are taken into consideration. The only disparate trend at odds with the $\mathrm{BCU}$ scores is that Fine-tuning Qwen models that have been instruction-tuned stand out by performing extra poorly with respect to $\mathrm{CCU}$ score. We hypothesise that this is a consequence of an increase in $P_{M}(t_{C}|Q)$ (i.e. prediction probability without context) from the fine-tuning, yielding less room for improvement in prediction confidence when context is introduced.
@@ A.2 Analysis of inflated CMT performance on CounterFact
The inflated performance on CounterFact, observed in Figures˜2 and 3, can potentially be explained by a suboptimal default prompt for CounterFact. Following previous work, the default prompt only contained the example to be completed, without any additional instructions or few-shot examples. For NQ and DRUID, the default prompt contained task instructions and few-shot examples. Furthermore, we observe how Prompting performs best on CounterFact on average, with a near perfect performance, indicating that a better default prompt may have neutralised any additional improvements from other CMTs. This raises the question of whether certain CMTs only address low context utilisation when caused by poor prompting, finding no leverage if the prompt already is adequate.
@@ A.4 Performance of Relevance Judgement
For the Multi-agent technique, we investigate whether instruction-tuned LMs are capable of identifying irrelevant context when explicitly prompted to do so. According to Table˜6, the Multi-agent approach demonstrates strong performance in detecting irrelevant contexts and in recognising gold contexts as relevant. Although it does not reliably maintain a closed-book response when directly generating responses (i.e. Regular), it can accurately detect irrelevance when equipped with an explicit relevance assessment setup.
@@ A.5 Features Impacting Context Utilisation
See Table˜3 for the correlation values between model features and context utilisation. See Table˜7 for the correlation values between input features and context utilisation.
@@
Table 7: Spearman’s $\rho$ between BCU and different input aspects. Correlation vallues for Regular or with an absolute value above 0.3 are shown. Correlation values with an absolute value below 0.3 are marked in gray. Significant correlation values (p-value < 0.05) are marked in bold. Results are measured across models.